Please help!
If (x) = 2x -4 and g(x) = x^2 -4x, find each value
1. f(3) -1
***** I got 2 for an answer
f(3)-1
f(3)-1=2(3)-4
6-4=2
answer =2
but... what's the -1 for then? do i do anything with it?

Answer by KMST(5328)

If I see "f(3) - 1" , I interpret that as find the value of the function for x=3 and then subtract 1.
So f(3)=2, as you found.
f(3)-1=2-1=1